OpenStudy (anonymous):

A ladder is resting against a wall. The ladder and the ground make an angle of 45° and the ladder is 7 ft from the wall. How long is the ladder?

OpenStudy (anonymous):

\[7\times \sqrt{2}\]

OpenStudy (anonymous):

9.90 feet?

OpenStudy (anonymous):

yeah 7/cos45=9.899=9.90 cos(x)=a/h h=a/cos(x)